OpenStudy (josephreak):

Cubic roots of 1 (complex numbers) Hi, How should I go on about solving the cubic roots of 1 or (omega) on my TI 89T. I looked it up and I couldn't find anything about it, I dunno if I'm using the correct terms or not. (I don't study in English) Anyways, here goes: (the sum) => w+w^2+1=0 (multiply) => w*w^2*1 = 1 How do I go on about solving this: (2+5w+2w^2)^2 By hand the correct answer (I hope lol) should be 9w^2 Thanks!

TheSmartOne (thesmartone):

Are you trying to expand it, factor it, or solving for w?

TheSmartOne (thesmartone):

I would suggest first expanding it, and then substitute u = w^2 to factor it.

OpenStudy (josephreak):

Hm, I dunno. I defined \[\omega = -\frac{ 1 }{ 2 }+\frac{ \sqrt{3} }{ 2 }i\] then I calculated it and the answer is correct, it was: \[(2+5\omega+2\omega^2)^2\] (sub w for what we defined) and the answer (on calc screen) is: \[-\frac{ 9 }{ 2 }-\frac{ 9\sqrt{3} }{ 2 }i\] divide by w^2 and you get 9. So it does calculate them when I define it into it, but it would be a lot better if I could make it display it symbolically instead of sub'ing. Anyway to do that?

OpenStudy (phi):

It's not clear (to me) what you are trying to answer. if the question is How should I go on about solving the cubic roots of 1 use the idea that 1 = exp( n 2 pi i) for n=1,2,3 with n=1, write 1 as exp( i 2 pi) take the cube root: exp(i ⅔ pi) convert exp( i A) to rectangular coords using cos(A) + i sin(A) do the same thing for n=2 and n=3 (n=4 will repeat these 3 solutions)

OpenStudy (josephreak):

Finding the cube roots of 1 isn't what we're looking for here. We know that cube root of 1 has 3 roots: \[1, -\frac{ 1 }{ 2 }+\frac{ \sqrt{3} }{ 2 }i,-\frac{ 1 }{ 2 }-\frac{ \sqrt{3} }{ 2 }i\] so we replace the two imaginary roots with a w, and 1. 1, w, w^2 (order) sum of all = 0 multiply of all = 0 and the question comes in the form of w equation, like the one I posted. I don't know how to explain it good enough tbh

OpenStudy (phi):

ok, but I assume you mean multiply of all = 1 ?

OpenStudy (josephreak):

yeah, sorry, my bad.

OpenStudy (phi):

so, you are asking, given the two equations 1+w+w^2 = 0 1*w*w^2 = 1 how do we find all 3 unique solutions ?

OpenStudy (josephreak):

Given those two equations, evaluate (2+5w+2w^2)^2

OpenStudy (josephreak):

The correct answer should be 9w^2 my original question was to how to evaluate this using TI 89 lol

OpenStudy (phi):

If you mean symbolically, no idea if you mean numerically, then the manual will help http://brownmath.com/ti83/complx89.htm

OpenStudy (josephreak):

---------------------------------- w+w^2+1=0 w*w^2*1=0 ---------------------------------- I'll put up an example for clarification. If 1, w and w^2 are the cubic roots of 1, prove that: \[(1+\omega^2)^8=\omega^2\] We know that 1+w^2 = -w substitute in the equation and evaluate (-w)^8 which is: w^8 which equals to w^2 I would like the TI 89 to do that, if possible.